What is the length of fencing material, in feet, needed to enclose this yard?. that is 35 feet and 20 feet.
Answer by fractalier(6550)

Essentially you are asking to find the perimeter of a rectangular yard.
Perimeter P = 2L + 2W = 2(35) + 2(20) = 70 + 40 = 110 feet of fencing